OpenStudy (anonymous):

−1.66=−5.09+2.17ln(x) |Could someone help me solve this Step by step please

OpenStudy (anonymous):

First organize your equation to read: −5.09+2.17ln(x) = -1.66 1) Add 5.09 to both sides 2.17ln(x) = -1.66 2) Divide by 2.17 on both sides ln(x) = -1.66/2.17

OpenStudy (anonymous):

@pabloecortez07 It says "non real calculation"....

OpenStudy (anonymous):

Im sorry, I did not add 5.09 to -1.66 So : 1) Add 5.09 to both sides 2.17ln(x) = 3.43 2) ln(x) = 3.43/2.17

OpenStudy (anonymous):

Thanks for da help @pabloecortez07

OpenStudy (agent0smith):

Now Put both sides to the power of e to find x.